OpenStudy (genny7):

Choose the two equations you would use to solve the absolute value equation below. Then solve the two equations. |2x + 5| = 3 A. 2x + 5 = 3 and 2x + 5 = –3; {–1, –4} B. 2x + 5 = 3 and 2x + 5 = –3; {1, –4} C. 2x – 5 = 3 and 2x + 5 = –3; {–1, 4} D. 2x – 5 = 3 and 2x + 5 = –3; {–1, –4}

OpenStudy (genny7):

@vishweshshrimali5

OpenStudy (vishweshshrimali5):

Just like this previous one: \[\large{|x+a| = b}\] \[\large{\implies x+a = b ,x+a = -b}\] I can extend it to this: \[\large{|cx + a| = b}\] \[\large{\implies cx+a = b, ~~cx+a = -b}\]

OpenStudy (vishweshshrimali5):

What do you think the answer should be ?

OpenStudy (genny7):

I think its B

OpenStudy (vishweshshrimali5):

Your equations are correct. Check the values of x you will get from them

OpenStudy (genny7):

ok, thanks your a really great help! :)

OpenStudy (vishweshshrimali5):

Your welcome. And the correct answer would be A

OpenStudy (genny7):

lol thanks :)